OpenStudy (anonymous):

Suppose the kick is attempted 36.0m from the goalposts, whose crossbar is 3.05 m above the ground. If the football is directed perfectly between the goalposts, will it pass over the bar and be a field goal? Show why or why not. If not, from what horizontal distance must this kick be made if it is to score? apply x=xo+voxt+1/2axt^2 x0=0, x=36.0m, vox=16.0m/s, ax=0m/s^2 solver for t apply y=y0+voyt+1/2ayt^2 y0=0, y=?, voy=12.0m/s, ay=-9.8m/s^2, t= previouslyfound solve for y i need step by step help please

OpenStudy (amistre64):

you have a quadratic in t .... hard to read the notation tho

OpenStudy (amistre64):

and they give you everything needed to substitute in to find y

OpenStudy (anonymous):

yes it is sorry if its hard to read

OpenStudy (amistre64):

how do we solve a quadratic equation? there are multiple methods available to us

OpenStudy (anonymous):

you need to find the times that y = 3.05 m and plug the later one into x(t) to see if it is greater than 36 m

OpenStudy (anonymous):

i can try and tell you what i geto for it . when id id itth e first time the numbers didn't look right

OpenStudy (amistre64):

x=xo+voxt+1/2axt^2 x0=0, x=36.0m, vox=16.0m/s, ax=0m/s^2 36=0+16t+1/2(0)t^2 hmm, if ax=0 then t is linear

OpenStudy (anonymous):

for x i got 2.2

OpenStudy (amistre64):

so we are given a linear equation and asked to solve for the time, the time is then applied to an equation that models time spent in the air 36 = 16t, t = 36/16 = 18/8 = 9/4

OpenStudy (anonymous):

solve for t with the y equation... you will get 2 solutions: one as the ball goes up and one as the ball comes down. you use the time for when the ball is coming down to see how far it has travelled.

OpenStudy (amistre64):

so it takes 2.25 seconds to travel 36 meters, y is equal to or greater then the height of the goal post, then we have scored

OpenStudy (amistre64):

y=y0+voyt+1/2ayt^2 y0=0, y=?, voy=12.0m/s, ay=-9.8m/s^2, t= previouslyfound solve for y y=0 +12t+1/2(-9.8)t^2 and t is known

OpenStudy (anonymous):

\[3.05 = 0 +12t \frac{m}{\sec} -\frac{1}{2}\cdot 9.8t^2 \frac{m}{\sec^2}\]solve for t.

OpenStudy (amistre64):

pgs approach will work, but its not 'according' to the problems instructions :)

OpenStudy (anonymous):

7.1

OpenStudy (anonymous):

thet're doing it but in a different way... you can solve for t using the x equation and then plug in that time to the y equation and see if the height is above or below the goalpost.

OpenStudy (anonymous):

\[36=0+16t \Rightarrow t=\frac{36}{16}=\frac{9}{4}\] so \[y \left( \frac{9}{4} \right)=0+12\left( \frac{9}{4} \right)-4.9\left( \frac{9}{4} \right)^2 = 2.19375 \text{ m}\]

OpenStudy (anonymous):

no good

OpenStudy (anonymous):

okay hold on ... i'mg ettingth at t=2.11

OpenStudy (anonymous):

to find out how far, solve for t using the y equation with y(t) = 3.05m. plug that into the x equation and solve for x to find out how far to get a field goal with those initial conditions.

OpenStudy (anonymous):

okay i will try and see

OpenStudy (anonymous):

30.5=0+16.0m/s+2.19375+1/2+0t^2

OpenStudy (anonymous):

okay so .163?

OpenStudy (anonymous):

? which part are you talking about?

OpenStudy (anonymous):

or 24.81 the part where youte ll me to plug it backi n

OpenStudy (anonymous):

if you solve for t using the y equation you get... \[3.05 = 0 +12t - 4.9t^2 \Rightarrow t = 0.28804 \text{ & } 2.16093\] \[x \left( 2.16093 \right)=16(2.16093) = 34.57488 \text{ m}\]

OpenStudy (anonymous):

okay let me retype it and see...

OpenStudy (anonymous):

okay yes i do get that number

OpenStudy (anonymous):

now do we subtract 34.57488 and 2.19375m=32.38m

OpenStudy (anonymous):

no! 34.57488 m is how far out in order to make the field goal. the 2.19375 m is the height of the ball when kicked from 36 m out.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):